OpenStudy (anonymous):

20.3 is 33 1/3 % of what number?

OpenStudy (anonymous):

60, roughly.

OpenStudy (anonymous):

33 1/3 = 1/3 so your number is 20.3 * 3 = 60.9

OpenStudy (anonymous):

or not.